# Compact manifolds of dimension $n\geq 12$ with positive isotropic curvature

### v21

@@ 1 Introduction
The notion of positive isotropic curvature was introduced by Micallef and Moore [63]. Recall that a Riemannian manifold $M$ (of dimension $\geq 4$) is said to have positive isotropic curvature (strictly PIC) if for all points $p\in M$ and all orthonormal 4-frames $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{p}M$, the curvature tensor satisfies
@@
(By [64] the converse of Theorem 1.1 is also true.) In fact, from the proof of Theorem 1.1 we know that a component appearing in the connected sum decomposition of $M$ in Theorem 1.1 which is diffeomorphic to the total space of an orbifiber bundle over $\mathbb{S}^{1}$ or $\mathcal{I}$ with generic fiber diffeomorphic to $\mathbb{S}^{n-1}$ is diffeomorphic to $\mathbb{R}P^{n}\sharp\mathbb{R}P^{n}$, or $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$, or $\mathbb{S}^{n-1}\tilde{\times}\mathbb{S}^{1}$, and is diffeomorphic to a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries. Here $\mathbb{S}^{n-1}\tilde{\times}\mathbb{S}^{1}$ is the (unique) nonorientable $\mathbb{S}^{n-1}$-bundle over $\mathbb{S}^{1}$ with structure group $O(n)$. So Theorem 1.1 extends Theorem 1.4 in [20]. (See also the remark after the proof of Theorem 1.1.) Compare also the Main Theorem in [29]; here I don’t know that if a $\mathbb{S}^{n-1}/\Gamma$-bundle ($|\Gamma|\geq 2$) over $\mathbb{S}^{1}$ or $\mathcal{I}$ admits a metric with positive isotropic curvature, whether or not the structure group can always be reduced to the isometry group Isom$(\mathbb{S}^{n-1}/\Gamma)$. (I conjecture that if a $\mathbb{S}^{n-1}/\Gamma$-bundle ($n\geq 5,|\Gamma|\geq 2$) over $\mathbb{S}^{1}$ or $\mathcal{I}$ admits a metric with positive isotropic curvature, the structure group can always be reduced to Isom$(\mathbb{S}^{n-1}/\Gamma)$.) On the other hand, Chen-Tang-Zhu predicted on p. 47 of [29] that a weaker version of Theorem 1.1 should hold once Hamilton’s curvature pinching estimates in [48] can be extended to higher dimensions. For the definition of connected sum of smooth manifolds see for example, [3], [25], [58], [66] and [80].
@@
As already observed in Section D in [48], in the general case, the Ricci flow surgeries on neck-pinch singularities may introduce orbifold singularities even if we start with a manifold. In the orbifold case, a priori, an $\varepsilon$-neck or $\varepsilon$-tube may not be universally noncollapsed; this may lead to difficulty in some contradiction-via-compactness argument. Chen-Tang-Zhu [29] dealt with this difficulty by using the technique of pulling back the Ricci flow solutions restricted to these domains via Hamilton’s canonical uniformizations (cf. Section C.2 in [48] and the appendix of our paper); see the proof of Propositions 4.2 and 4.4 in [29]. Of course we can also use arbitrary universal covering maps to pull back the solutions restricted to these domains. Now by using Hamilton’s compactness theorem we are in the following typical situation: We have a strong $\varepsilon^{\prime}$-neck upstairs, which is located at some covering space and which may not be equivariant under the action of the deck transformation groups (recalling that the definition of Cheeger-Gromov convergence involves diffeomorphisms), but we want to get some strong $\varepsilon$-neck downstairs. Of course we may use Hamilton’s canonical uniformization in the final time slice, which is usually equivariant with respect to the action of the deck transformation groups and descends to give an $\tilde{\varepsilon}$-neck downstairs in the final time slice. But note that a Hamilton’s canonical parametrization in the final time slice may not be such one in the previous time slices. Does Hamilton’s canonical parametrization in the final time slice give a strong $\varepsilon$-neck downstairs? By using the argument in Section 2 of [46] one sees that it does. (This little detail was not written down in [29]. By the way, the word “strong” is missing in the statement and proof of Propositions 4.2 and 4.4 in [29]; moreover the points $x_{j}$ there may not be in the same horn $H$.) More generally, by using the argument in Section 2 of [46] and Kotschwar’s backwards uniqueness theorem for the Ricci flow [59] we show a neck strengthening lemma (see Lemma 5.8), which extends and improves Lemma 4.3.5 in [6]. Lemma 5.8 is used in the proof of Propositions 2.9, 2.10, 3.4 and 3.9. We also refer the reader to p. 101 of Bessières et al. [6] for some details for the application of their Lemma 4.3.5 in the construction of Ricci flow with surgery.
@@
The proof of our theorems depends crucially on Brendle’s work [20], and I was also inspired by Chen-Tang-Zhu’s work [29], the works of Besson and his collaborators [6] [7], and the work of Kleiner-Lott [56], etc. As mentioned above, we extensively use results and techniques from topology, for example, Cerf-Palais theorem, isotopy extension theorem and covering space technique. I also make great efforts to add some necessary details involved in the construction of Ricci flow with surgery on compact orbifolds, which were neglected in the previous related works. For examples, see the proof of Propositions 2.10, 3.4 and 5.4, and Lemmas 5.5 and 5.8.
@@
After this manuscript was submitted, we continue to investigate to improve its main results. Note that the Remark on p. 312 in Bredon [15] gives an equivariant ambient tubular neighborhood theorem, and more generally one can also derive an orbifold version of the ambient tubular neighborhood theorem; similarly there is an equivariant/orbifold version of the ambient collar theorem. (Using the latter we can get an orbifold version of Theorem 2.3 in Chapter 8 of [49].) We can use the ambient isotopy uniqueness of closed tubular neighborhoods of compact suborbifolds to prove fine properties of orbifold ancient $\kappa$-solutions and standard solutions when they contain caps with ends of type $\mathbb{S}^{n-1}/\Gamma\times(-1,1)$ with $\Gamma$ nontrivial, extending Propositions 2.5, 2.6 and 2.10, and strengthen the definition of $\varepsilon$-caps in Section 2 in these cases. Using this, we can show that in fact the weak connected sum in Theorem 1.3 can be strengthened to be connected sum, and furthermore, we can improve Theorem 1.1 to a more precise classification result; see [52]. Moreover, it is possible to extend the main results in this paper and [52] to the case of open manifolds of dimension $n\geq 12$ with uniformly positive isotropic curvature and bounded geometry; see [53].
@@
In Section 2 we classify odd ($\geq 3$)-dimensional spherical orbifolds with nonempty and isolated singularity set (for the even ($\geq 4$)-dimensional case see [29]), define $\varepsilon$-caps, and describe the canonical neighborhood structure of orbifold ancient $\kappa$-solutions and orbifold standard solutions. In Section 3 we prove the so called bounded curvature at bounded distance property, choose the cutoff parameters for surgical Ricci flow under the canonical neighborhood assumption, and construct an $(r,\delta)$-surgical solution to the Ricci flow starting with a compact, connected Riemannian orbifold of dimension $n\geq 12$ with isolated singularities and with positive isotropic curvature. In Section 4, we prove Theorem 1.3 and Theorem 1.1 using the construction in Section 3, and give a proof of Corollary 1.2 using Theorem 1.1. In Section 5, which is an appendix, we introduce various notions on necks and (topological) caps, and establish results on the composition of $\varepsilon$-isometries and the gluing and strengthening of $\varepsilon$-necks, which are used in the previous sections.
@@ 2 Orbifold ancient $\kappa$-solutions and standard solutions
Let $(\mathcal{O},g)$ be a connected Riemannian orbifold. Following [56], we call a path in $\mathcal{O}$ as defined in Section 2.2.1 of [9] a special curve. We emphasize that for a special curve $\gamma:I\rightarrow\mathcal{O}$, to each $t\in I$ (here $I\subset\mathbb{R}$ is an interval) such that $|\gamma|(t)$ is a singular point, there is associated a smooth local lift of $\gamma$ around $t$ which is given as part of the data. (See also Definition 3.1 in [14].) Given a special curve $\gamma$ in $\mathcal{O}$, we can pullback the tangent bundle $T\mathcal{O}$ and the Levi-Civita connection of $(\mathcal{O},g)$ via $\gamma$, and define parallel transport along $\gamma$ as in the case of Riemannian manifolds. (One can also define piecewise smooth special curves and parallel transport along a piecewise smooth special curve.) We also define homotopy of special curves as in Section 2.2.1 of [9] and Section 2.2 of [56]. (We can also define homotopy of piecewise smooth special curves by using piecewise smooth homotopies in local models; for the manifold case see for example Appendix 7 in [57].)
@@
As in [18], let $\mathcal{C}_{B}(\mathbb{R}^{n})$ be the vector space of algebraic curvature tensors on $\mathbb{R}^{n}$. Recall ([23][18]) that the curvature tensor of a $n$-dimensional Riemannian orbifold $(\mathcal{O},g)$ lies in the PIC1 cone in $\mathcal{C}_{B}(\mathbb{R}^{n})$ (see item (ii) of Definition 1.1 in [20]) if and only if the curvature tensor of $\mathcal{O}\times\mathbb{R}$ lies in the PIC cone (the set of all algebraic curvature tensors with nonnegative isotropic curvature) in $\mathcal{C}_{B}(\mathbb{R}^{n+1})$, and the curvature tensor of $(\mathcal{O},g)$ lies in the PIC2 cone in $\mathcal{C}_{B}(\mathbb{R}^{n})$ (see item (iii) of Definition 1.1 in [20]) if and only if the curvature tensor of $\mathcal{O}\times\mathbb{R}^{2}$ lies in the PIC cone in $\mathcal{C}_{B}(\mathbb{R}^{n+2})$. (Here, at a singular point, we should look at the curvature tensor at the corresponding point in a local model.) Recall that the curvature tensor of $(\mathcal{O},g)$ lies in the PIC2 cone if and only if $(\mathcal{O},g)$ has nonnegative complex sectional curvature; a proof (attributed to N. Wallach) of this equivalence appeared in a 2007 arXiv preprint of L. Ni and J. Wolfson, see also for example Proposition 7.18 in [18]. When (at each point) the curvature tensor of $(\mathcal{O},g)$ lies in the interior of the PIC2 cone, we also say that $(\mathcal{O},g)$ is strictly PIC2 (in this case, the inequality in item (iii) of Definition 1.1 in [20] is strict; see also Section 2 in [22]); this condition is equivalent to that $(\mathcal{O},g)$ has positive complex sectional curvature. (One can also consult the first paragraph on p. 55 and Chapter 7 of [18].)
@@
Let $(\mathcal{O},g(t)),t\in[0,T]$, be a complete solution to the Ricci flow. Suppose that $(\mathcal{O},g(t)),t\in[0,T]$, is weakly PIC2 and strictly PIC, and there is a point $x_{0}$ such that the Ricci curvature of $g(0)$ at $x_{0}$ is positive. Then for any $t>0$ and any $x\in\mathcal{O}$, the curvature tensor of $g(t)$ at $x$ lies in the interior of the PIC2 cone.
Proof.   Otherwise, there is a space-time point $(p,t^{\prime})$ with $t^{\prime}>0$ such that the curvature tensor at $(p,t^{\prime})$ lies on the boundary of the PIC2 cone. By the orbifold version of Proposition 6.6 in [20] for all $0<t\leq t^{\prime}$ the universal cover of $(\mathcal{O},g(t))$ splits off a line. In particular, the Ricci curvature of $(\mathcal{O},g(t))$ cannot be strictly positive for $0<t\leq t^{\prime}$. On the other hand, by assumption the Ricci curvature of $g(0)$ is positive at the point $x_{0}$, so the Ricci curvature is positive everywhere for any $t>0$ by the strong maximum principle for systems (Proposition 12.47 in [35]); here note that the reaction term in the evolution equation for the Ricci tensor is weakly positive definite under the weakly PIC2 assumption. A contradiction. $\Box$
@@
(cf. Corollary 6.7 in [20], and Theorem 3.4 in [29])    Let $n\geq 5$, and $(\mathcal{O},g(t))$, $t\in(-\infty,T]$, be an orbifold ancient $\kappa$-solution of dimension $n$ (with at most isolated singularities). Suppose that $(\mathcal{O},g(t))$ satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in PIC$ for some uniform constant $\theta>0$, and there is a spacetime point $(x_{0},t_{0})$ such that the curvature tensor at $(x_{0},t_{0})$ lies on the boundary of the PIC2 cone. Then for each $t$, $(\mathcal{O},g(t))$ is isometric to a shrinking Ricci soliton $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ or $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}\mathbb{R}$ for some finite subgroup $\Gamma$ of $O(n)$ acting freely on $\mathbb{S}^{n-1}$. In particular, if $\mathcal{O}$ has no singularities and has exactly one end, it must be diffeomorphic to some $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus\bar{B}$, where $\sigma$ is an isometric involution of the spherical space form $\mathbb{S}^{n-1}/\Gamma$ with no fixed points, and $\hat{\sigma}$ and $B$ are defined as in the appendix; if $\mathcal{O}$ has nonempty isolated singularities, it must be diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus\bar{B}$ as defined in the appendix. Furthermore, there exists a positive constant $\varepsilon_{4}=\varepsilon_{4}(n)\leq\varepsilon_{1}$, where $\varepsilon_{1}$ is the constant in Proposition 5.4, with the following property.
@@
Proof.   Compare the proof of Corollary 6.7 and Proposition 6.17 in [20] and Theorem 3.4 in [29]. We pull back our solution to the universal cover and get $(\widetilde{\mathcal{O}},\tilde{g}(t))$. As noted above the orbifold version of Proposition 6.6 in [20] holds. It follows that at any time $t$, $(\widetilde{\mathcal{O}},\tilde{g}(t))$ is isometric to a product $(X,g_{X}(t))\times\mathbb{R}$, where $X$ is a smooth manifold. Since by assumption $(X,g_{X}(t))\times\mathbb{R}$ is nonflat, weakly PIC2 and satisfies $\text{Rm}-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in\text{PIC}$ for some uniform constant $\theta>0$, the ancient solution $(X,g_{X}(t))$ is weakly PIC2 and uniformly PIC1. By Theorem 6.4 in [20], we see that at any time $t\leq t_{0}$, $(X,g_{X}(t))$ is a round $(n-1)$-sphere. Now we see that $(\mathcal{O},g(t))$ is a metric quotient of the evolving round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries (compare Theorem 13.3.10 in [74]). Then it follows from the $\kappa$-noncollapsing assumption on all scales that $\mathcal{O}$ is noncompact (cf. p.212 in [31] and p.52 in [29]). So $\mathcal{O}$ has exactly one or two ends, moreover, if $\mathcal{O}$ has two ends, it must split metrically as $\mathcal{O}^{\prime}\times\mathbb{R}$, where $\mathcal{O}^{\prime}$ is a metric quotient of $\mathbb{S}^{n-1}$; see Corollary 3.3 in [56]. As $\mathcal{O}$ has at most isolated singularities, if it has two ends, it must be isometric to $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ for some finite subgroup $\Gamma$ of $O(n)$ acting freely on $\mathbb{S}^{n-1}$, in particular, in this case $\mathcal{O}$ is a smooth manifold.

### v22

@@ 1 Introduction
The notion of positive isotropic curvature was introduced by Micallef and Moore [66]. Recall that a Riemannian manifold $M$ (of dimension $\geq 4$) is said to have positive isotropic curvature ((strictly) PIC) if for all points $p\in M$ and all orthonormal 4-frames $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{p}M$, the curvature tensor satisfies
@@
(By [67] the converse of Theorem 1.1 is also true.) In fact, from the proof of Theorem 1.1 we know that a component appearing in the connected sum decomposition of $M$ in Theorem 1.1 which is diffeomorphic to the total space of an orbifiber bundle over $\mathbb{S}^{1}$ or $\mathcal{I}$ with generic fiber diffeomorphic to $\mathbb{S}^{n-1}$ is diffeomorphic to $\mathbb{R}P^{n}\sharp\mathbb{R}P^{n}$, or $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$, or $\mathbb{S}^{n-1}\tilde{\times}\mathbb{S}^{1}$, and is diffeomorphic to a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries. Here $\mathbb{S}^{n-1}\tilde{\times}\mathbb{S}^{1}$ is the (unique) nonorientable $\mathbb{S}^{n-1}$-bundle over $\mathbb{S}^{1}$ with structure group $O(n)$. So Theorem 1.1 extends Theorem 1.4 in [21]. (See also the remark after the proof of Theorem 1.1.) Compare also the Main Theorem in [31]; here I don’t know that if a $\mathbb{S}^{n-1}/\Gamma$-bundle ($|\Gamma|\geq 2$) over $\mathbb{S}^{1}$ or $\mathcal{I}$ admits a metric with positive isotropic curvature, whether or not the structure group can always be reduced to the isometry group Isom$(\mathbb{S}^{n-1}/\Gamma)$. (I conjecture that if a $\mathbb{S}^{n-1}/\Gamma$-bundle ($n\geq 5,|\Gamma|\geq 2$) over $\mathbb{S}^{1}$ or $\mathcal{I}$ admits a metric with positive isotropic curvature, the structure group can always be reduced to Isom$(\mathbb{S}^{n-1}/\Gamma)$.) On the other hand, Chen-Tang-Zhu predicted on p. 47 of [31] that a weaker version of Theorem 1.1 should hold once Hamilton’s curvature pinching estimates in [51] can be extended to higher dimensions. For the definition of connected sum of smooth manifolds see for example, [3], [26], [61], [69], [71] and [86].
@@
As already observed in Section D in [51], in the general case, the Ricci flow surgeries on neck-pinch singularities may introduce orbifold singularities even if we start with a manifold. In the orbifold case, a priori, an $\varepsilon$-neck or $\varepsilon$-tube may not be universally noncollapsed; this may lead to difficulty in some contradiction-via-compactness argument. Chen-Tang-Zhu [31] dealt with this difficulty by using the technique of pulling back the Ricci flow solutions restricted to these domains via Hamilton’s canonical uniformizations (cf. Section C.2 in [51] and the appendix of our paper); see the proof of Propositions 4.2 and 4.4 in [31]. Of course we can also use arbitrary (smooth) universal covering maps to pull back the solutions restricted to these domains. Now by using Hamilton’s compactness theorem we are in the following typical situation: We have a strong $\varepsilon^{\prime}$-neck upstairs, which is located at some covering space and which may not be equivariant under the action of the deck transformation groups (recalling that the definition of Cheeger-Gromov convergence involves diffeomorphisms), but we want to get some strong $\varepsilon$-neck downstairs. Of course we may use Hamilton’s canonical uniformization in the final time slice, which is usually equivariant with respect to the action of the deck transformation groups and descends to give an $\tilde{\varepsilon}$-neck downstairs in the final time slice. But note that a Hamilton’s canonical parametrization in the final time slice may not be such one in the previous time slices. Does Hamilton’s canonical parametrization in the final time slice give a strong $\varepsilon$-neck downstairs? By using the argument in Section 2 of [49] one sees that it does. (This little detail was not written down in [31]. By the way, the word “strong” is missing in the statement and proof of Propositions 4.2 and 4.4 in [31]; moreover the points $x_{j}$ there may not be in the same horn $H$.) More generally, by using the argument in Section 2 of [49] and Kotschwar’s backwards uniqueness theorem for the Ricci flow [62] we show a neck strengthening lemma (see Lemma A.8), which extends and improves Lemma 4.3.5 in [6]. Lemma A.8 is used in the proof of Propositions 2.9, 2.10, 3.4 and 3.9. We also refer the reader to p. 101 of Bessières et al. [6] for some details for the application of their Lemma 4.3.5 in the construction of Ricci flow with surgery.
@@
The proof of our theorems depends crucially on Brendle’s work [21], and I was also inspired by Chen-Tang-Zhu’s work [31], the works of Besson and his collaborators [6] [7], and the work of Kleiner-Lott [59], etc. As mentioned above, we extensively use results and techniques from topology, for example, Cerf-Palais theorem, isotopy extension theorem and covering space technique. I also make great efforts to add some necessary details involved in the construction of Ricci flow with surgery on compact orbifolds, which were neglected in the previous related works. For examples, see the proof of Propositions 2.10, 3.4 and A.4, and Lemmas A.5 and A.8.
@@
After this manuscript was submitted, we continue to investigate to improve its main results. Note that the Remark on p. 312 in Bredon [16] gives an equivariant ambient tubular neighborhood theorem, and more generally one can also derive an orbifold version of the ambient tubular neighborhood theorem; similarly there is an equivariant/orbifold version of the ambient collar theorem. (Using the latter we can get an orbifold version of Theorem 2.3 in Chapter 8 of [52].) We can use the ambient isotopy uniqueness of closed tubular neighborhoods of compact suborbifolds to prove fine properties of orbifold ancient $\kappa$-solutions and standard solutions when they contain caps with ends of type $\mathbb{S}^{n-1}/\Gamma\times(-1,1)$ with $\Gamma$ nontrivial, extending Propositions 2.5, 2.6 and 2.10, and strengthen the definition of $\varepsilon$-caps in Section 2 in these cases. Using this, we can show that in fact the weak connected sum in Theorem 1.3 can be strengthened to be connected sum, and furthermore, we can improve Theorem 1.1 to a more precise classification result; see [55]. Moreover, it is possible to extend the main results in this paper and [55] to the case of open manifolds of dimension $n\geq 12$ with uniformly positive isotropic curvature and bounded geometry; see [56]. Recently, Zhengnan Chen [35] extended Brendle’s curvature pinching estimates to the case of $n\geq 9$. So combined with Chen’s estimates, our arguments imply that all these results also hold true for $9\leq n\leq 11$.
@@
In Section 2 we classify odd ($\geq 3$)-dimensional spherical orbifolds with nonempty and isolated singularity set (for the even ($\geq 4$)-dimensional case see [31]), define $\varepsilon$-caps, and describe the canonical neighborhood structure of orbifold ancient $\kappa$-solutions and orbifold standard solutions. In Section 3 we prove the so called bounded curvature at bounded distance property, choose the cutoff parameters for surgical Ricci flow under the canonical neighborhood assumption, and construct an $(r,\delta)$-surgical solution to the Ricci flow starting with a compact, connected Riemannian orbifold of dimension $n\geq 12$ with isolated singularities and with positive isotropic curvature. In Section 4, we prove Theorem 1.3 and Theorem 1.1 using the construction in Section 3, and give a proof of Corollary 1.2 using Theorem 1.1. In the appendix, we introduce various notions on necks and (topological) caps, construct orbifold standard solutions, and establish results on the composition of $\varepsilon$-isometries and the gluing and strengthening of $\varepsilon$-necks, which are used in the previous sections.
@@ 2 Orbifold ancient $\kappa$-solutions and standard solutions
Let $(\mathcal{O},g)$ be a connected Riemannian orbifold. Following [59], we call a path in $\mathcal{O}$ as defined in Section 2.2.1 of [9] a special curve. We emphasize that for a special curve $\gamma:I\rightarrow\mathcal{O}$, to each $t\in I$ (here $I\subset\mathbb{R}$ is an interval) such that $|\gamma|(t)$ is a singular point, there is associated a smooth local lift of $\gamma$ around $t$ which is given as part of the data. (See also Definition 3.1 in [14] or Definition 2.1 in [15].) Given a special curve $\gamma$ in $\mathcal{O}$, we can pullback the tangent bundle $T\mathcal{O}$ and the Levi-Civita connection of $(\mathcal{O},g)$ via $\gamma$ (cf. [34] and Section 6 of [15]), and define parallel transport along $\gamma$ as in the case of Riemannian manifolds. (One can also define piecewise smooth special curves and parallel transport along a piecewise smooth special curve.) We also define homotopy of special curves as in Section 2.2.1 of [9] and Section 2.2 of [59]. (We can also define homotopy of piecewise smooth special curves by using piecewise smooth homotopies in local models; for the manifold case see for example Appendix 7 in [60].)
@@
As in [19], let $\mathcal{C}_{B}(\mathbb{R}^{n})$ be the vector space of algebraic curvature tensors on $\mathbb{R}^{n}$. Following the notation in [85], let $C_{\text{PIC}}$, $C_{\text{PIC}1}$ and $C_{\text{PIC}2}$ be the subsets of $\mathcal{C}_{B}(\mathbb{R}^{n})$ which consist of algebraic curvature tensors of nonnegative isotropic curvature (weakly PIC, see item (i) of Definition 1.1 in [21]), of weakly PIC1 (see item (ii) of Definition 1.1 in [21]), and of weakly PIC2 (see item (iii) of Definition 1.1 in [21]) respectively. Recall ([24][19]) that the curvature tensor of an $n$-dimensional Riemannian orbifold $(\mathcal{O},g)$ lies in the cone $C_{\text{PIC}1}$ if and only if at each point the Riemannian orbifold $\mathcal{O}\times\mathbb{R}$ with the product metric has nonnegative isotropic curvature, and the curvature tensor of $(\mathcal{O},g)$ lies in the cone $C_{\text{PIC}2}$ if and only if at each point the Riemannian orbifold $\mathcal{O}\times\mathbb{R}^{2}$ with the product metric has nonnegative isotropic curvature. (Here, at a singular point, we should look at the curvature tensor at the corresponding point in a local model.) Note that the curvature tensor of $(\mathcal{O},g)$ lies in the cone $C_{\text{PIC}2}$ if and only if $(\mathcal{O},g)$ has nonnegative complex sectional curvature; a proof (attributed to N. Wallach) of this equivalence appeared in a 2007 arXiv preprint of L. Ni and J. Wolfson, see also for example Proposition 7.18 in [19]. When (at each point) the curvature tensor of $(\mathcal{O},g)$ lies in the interior of the cone $C_{\text{PIC}2}$, we also say that $(\mathcal{O},g)$ is strictly PIC2 (in this case, the inequality in item (iii) of Definition 1.1 in [21] is strict; see also Section 2 in [23]); this condition is equivalent to that $(\mathcal{O},g)$ has positive complex sectional curvature. (One can also consult the first paragraph on p. 55 and Chapter 7 of [19].)
@@
Let $(\mathcal{O},g(t)),t\in[0,T]$, be a complete solution to the Ricci flow. Suppose that $(\mathcal{O},g(t)),t\in[0,T]$, is weakly PIC2 and strictly PIC, and there is a point $x_{0}$ such that the Ricci curvature of $g(0)$ at $x_{0}$ is positive. Then for any $t>0$ and any $x\in\mathcal{O}$, the curvature tensor of $g(t)$ at $x$ lies in the interior of the cone $C_{\text{PIC}2}$.
Proof.   Otherwise, there is a space-time point $(p,t^{\prime})$ with $t^{\prime}>0$ such that the curvature tensor at $(p,t^{\prime})$ lies on the boundary of the cone $C_{\text{PIC}2}$. By the orbifold version of Proposition 6.6 in [21] for all $0<t\leq t^{\prime}$ the universal cover of $(\mathcal{O},g(t))$ splits off a line. In particular, the Ricci curvature of $(\mathcal{O},g(t))$ cannot be strictly positive for $0<t\leq t^{\prime}$. On the other hand, by assumption the Ricci curvature of $g(0)$ is positive at the point $x_{0}$, so the Ricci curvature is positive everywhere for any $t>0$ by the strong maximum principle for systems (Proposition 12.47 in [38]); here note that the reaction term in the evolution equation for the Ricci tensor is weakly positive definite under the weakly PIC2 assumption. A contradiction. $\Box$
@@
(cf. Corollary 6.7 in [21], and Theorem 3.4 in [31])    Let $n\geq 5$, and $(\mathcal{O},g(t))$, $t\in(-\infty,T]$, be an orbifold ancient $\kappa$-solution of dimension $n$ (with at most isolated singularities). Suppose that $(\mathcal{O},g(t))$ satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}$ for some uniform constant $\theta>0$, and there is a spacetime point $(x_{0},t_{0})$ such that the curvature tensor at $(x_{0},t_{0})$ lies on the boundary of the cone $C_{\text{PIC}2}$. Then for each $t$, $(\mathcal{O},g(t))$ is isometric to a shrinking Ricci soliton $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ or $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}\mathbb{R}$ for some finite subgroup $\Gamma$ of $O(n)$ acting freely on $\mathbb{S}^{n-1}$. In particular, if $\mathcal{O}$ has no singularities and has exactly one end, it must be diffeomorphic to some $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus\bar{B}$, where $\sigma$ is an isometric involution of the spherical space form $\mathbb{S}^{n-1}/\Gamma$ with no fixed points, and $\hat{\sigma}$ and $B$ are defined as in the appendix; if $\mathcal{O}$ has nonempty isolated singularities, it must be diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus\bar{B}$ as defined in the appendix. Furthermore, there exists a positive constant $\varepsilon_{4}=\varepsilon_{4}(n)\leq\varepsilon_{1}$, where $\varepsilon_{1}$ is the constant in Proposition A.4, with the following property.
@@
Proof.   Compare the proof of Corollary 6.7 and Proposition 6.17 in [21] and Theorem 3.4 in [31]. We pull back our solution to the universal cover and get $(\widetilde{\mathcal{O}},\tilde{g}(t))$. As noted above the orbifold version of Proposition 6.6 in [21] holds. It follows that at any time $t$, $(\widetilde{\mathcal{O}},\tilde{g}(t))$ is isometric to a product $(X,g_{X}(t))\times\mathbb{R}$, where $X$ is a smooth manifold. Since by assumption $(X,g_{X}(t))\times\mathbb{R}$ is nonflat, weakly PIC2 and satisfies $\text{Rm}-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}$ for some uniform constant $\theta>0$, the ancient solution $(X,g_{X}(t))$ is weakly PIC2 and uniformly PIC1. By Theorem 6.4 in [21], we see that at any time $t\leq t_{0}$, $(X,g_{X}(t))$ is a round $(n-1)$-sphere. Now we see that $(\mathcal{O},g(t))$ is a metric quotient of the evolving round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries (compare Theorem 13.3.10 in [78]). Then it follows from the $\kappa$-noncollapsing assumption on all scales that $\mathcal{O}$ is noncompact (cf. p.212 in [33] and p.52 in [31]). So $\mathcal{O}$ has exactly one or two ends, moreover, if $\mathcal{O}$ has two ends, it must split metrically as $\mathcal{O}^{\prime}\times\mathbb{R}$, where $\mathcal{O}^{\prime}$ is a metric quotient of $\mathbb{S}^{n-1}$; see Corollary 3.3 in [59]. As $\mathcal{O}$ has at most isolated singularities, if it has two ends, it must be isometric to $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ for some finite subgroup $\Gamma$ of $O(n)$ acting freely on $\mathbb{S}^{n-1}$, in particular, in this case $\mathcal{O}$ is a smooth manifold.